OpenStudy (anonymous):

The sum of the measure of the interior angles of a polygon is 1080°. What polygon is it? hexagon triangle decagon octagon

hero (hero):

Basically test each polygon (except for the triangle). We know triangle definitely isn't it: Test Hexagon 360/6 = 60 (exterior angle) 180 - 60 = 120 (interior angle) 120*6 = 720 (Sum of interior) Test Decagon: 360/10 = 36 (exterior angle) 180 - 36 = 144 (interior angle) 144 x 10 = 1440 (Sum of Interior) Test Octagon: 360/8 = 45 (exterior angle) 180 - 45 = 135 (interior angle) 135*8 = 1080 (Sum of Interior) Therefore, the polygon with a sum of interior angles of 1080 degrees is an octagon.